SEPARATING TRANSFORMATIONS FOR CONTROLLED VARIABLE MOLECULAR DYNAMICS

BEN LEIMKUHLER

Abstract. We show that Nosé’s thermostat can be given a mechanical formulation (with Hamiltonian having a Euclidean kinetic energy metric). We also give a mechanical formulation for simultaneous control of temperature and pressure. This approach simplifies the construction of symplectic methods while providing a more intuitive perspective on the nature of the controlled variable molecular dynamics.